AO Commissioning on 13 October 2011

Observer(s):C. Arcidiacono, L. Busoni, A. Puglisi, JC. Guerra, K. Boutsia, J. Brynnel (LBTO)

Telescope Operator: D. Huerta (LBTO)

SW Support: T. Sargent (Tucson)

Telescope Support: M. Midkiff (LBTO)

Instrument Support: J. Morris (LBTO)

Summary

Clear sky when we open and winds <5m/s. There are problems with the initial collimation because part of the pupil is vignetted.
Check observer routines - Data directory:/OldRepository/PISCES_Data/20111013/
We had problems with the telescope. In particular, we lost ~20min because of power failure in the rack, ~20min because we could not abort a preset, ~1hr because the rotator was in fault mode
The AO group went on with their tests (for details see Lorenzo's email). At the end of the night we obtained 4 standard stars in different elevations in 3 filtes (JHK).
At dawn we took skyflats in J and H and a set of darks for calculating corquad coefficients, but the lights were on.

Details

01:23 Start opening procedure

01:38 Checking pisces status we find it busy - 600sec integration
  • while they were taking some test darks, the filter has been changed from dark to FeII.

01:44 We have the same collimation problems as yesterday.

01:55 preset, OBJECT=BS9188, MODE=active, GS=13, RS=0, PA=-95 - with the observe routine
  • there is vignetting to the pupil in the wfs.

02:37 preset to the standard BS9188 to test offsets in open loop (MODE=active)
  • preset failed giving error:
% Execution halted at: TRIGGER_WFSDX       5
  /home/LBTO/idl/wfsc/observe/trigger_wfsdx.pro
  • it exits again because: => Exiting the script because the loop opened
  • when we do iif_offset for set_position we give: iif_offset, -5,-5, COORD=DETXY but IDL tells us that it has executed: => OffsetPointing 0 -5 -5 RADEC ABS right
  • we need to be able to exit the script before finishing it in case we need to change something....
  • we send the preset again increasing the exptime ....
  • with 30sec we don't see the star - so we give up on the standard and the AO guys start their test.

02:56 Preset again on BS9188 in ACE mode...

02:58 We don't see any stars on Pisces because there is STILL the stop plate in front of the camera.

03:04 The plate is off. Preset again in mode ACE.
  • preset failed because no guide star found.
  • check pointing
  • repeat preset - failed
  • checking pointing again, because we cannot find the guide star.

03:24 preset again to Astar_45 (Rmag=8.7)
  • preset, OBJECT=Astar_45 MODE=ace, GS=2, RS=0, PA=-103
  • off mirror ventilation
  • failed giving error at the outgain: "IDL variable undefined 'gainv'"

03:49 AO problem seems resolved - we send the preset again.

TEST DITHER SCRIPT.

Astar_45dither.obs To execute a dither pattern in the position (0,0) is necessary to run the iif_offset x,y, coord='detxy' from another IDl terminal with the values x and y equal to the offset applied to center the star in position 0,0
It would be useful, when the dither pattern finishes, to bring the star back to the central position, without taking an image.

Orientation RADEC of PISCES.

Movements in RADEC to determine the North/East orinetation. Two images taken: RA=0,DEC=2 , image=pisces.20111013.0054.fits Ra=2,DDEC=0, Image=pisces.20111013.0055.fits.

Going on with the AO tests

04:28 Check if AO can point low - we preset to Astar_23 (elev 35)
  • preset, OBJECT=Astar_45 MODE=ace, GS=5, RS=0, PA=-72
  • preset cancelled because of hardware problem. W switched off..

04:43 Telescope is loosing power - we need to switch everything off!
  • we lost power in the whole rack.

05:06 problem solved - start with the presets.

05:20 Preset sent to test a preset close loop at low elevation. The preset failed by problems with the pupil center procedure in the W-unit. After analysis of the error, the new preset was sent.

5:25 Preset Ok

5:39 telescope to zenith to check power failure.

6:04 Preset. Astar_189 (R=11.7). preset,Astar_189,mode=ace,gs=11,rs=0,pa=90 Failed by CCD47 not ready. Message from Observe: preset inactive. The preset is resend. Failed. Message from Observe: preserve inactive. The command preset was sent to the telescope and it wasn't arrived to the WFS.

6:11 Resend the preset. Ok.
  • taking sequence of images. The AO loop is open due to problem with the rerotator (idle), this probably is problem of the message deamon.

06:34 checking autogain. Present sent to telescope and to the AO, the Observe finished and acquire an image when the AO system was still doing the autogain. The AOS showed that the loop was closed. Failed.

06:54 Preset. ok. Send another preset to test a modified version of the autogain procedure to report error if the procedure doesn't success. The Observe routine shows: error........IDL_SAFE_SKIP_SAFE. The AOS reports the same error. The AO close loop procedure stops.

7:14 Preset to chi_persei Sending the Observe gave an error of multiple catalogue. The catalogue was saved in two different directories: /home/lbto/idl/catalogs/ /home/lbto/pisces/

The preset is sent failed during reading the catalogue, but the script send the preset to the AO and close the loop using the previous star preset.

The catalogue is checked and it was found a problem during edit

7:44 preset failed - active optics loop has an error
  • the wavefront sensor is saturated - need to change guide star
  • preset failed again because GS is seen in guider but not from AO.
  • repeat preset
  • with FeII we have very few counts for this star and we change to H.

8:09 DIMM - 0.9''

8:10 Telescope zenith ND filter out.

8:25 Preset to chi_persei. Field curvature analysis

08:30 trying to abort the preset from the telescope and it doesn't do it...

08:47 telescope problem solved - send preset chi_persei

08:51 not sure if it took the correct object - change filter back to FeII.

09:12 preset failed - no star found by AO. The guider though sees the star
  • repeat preset

09:43 preset failed because of rotator error (it is in fault mode)
  • this happened while they were trying to run on OB with the GS at 70'' of distance.
  • this test is to explore how far the GS can be.
  • to fix the rotator the TO needs to go up to the dome.

Standard star

10:35 preset for BS9107 (elev75) - standard star
  • first set filter Ks
  • preset failed giving error: ===> Active Optics loop is not running
  • the T.O has to correct pointing before repeating preset
image set starts:pisces.20111013.0187.fits 10:49 repeat preset BS9107 (el 72)
  • it took 5 iterations to collimate

11:04 change filter from observe routine (J) and repeat dither pattern. Launch the Observe with changing filter, and the Observe routine didn't change the filter and start to integrate. The filter was changed manually. image set starts:pisces.20111013.0195.fits 11:19 change to filter H. image: pisces.20111013.0201.fits. Change the filter outside of scripts.

11:23 preset for BS9118.(el 54). Filter H. Image set starts:pisces.20111013.0207.fits Failed... Error message: Active optics loop is not running. Resend the preset. Error message: Failed. Active optics loop is not running. Resend the preset(11:32). Ok

11:50 we tried to repeat the dither pattern with a different filter in the same script but it gave us this error
===> Error in irc_offsetpointing
        err = Resmsg[0]  = OffsetPointing result status: Error
        err = Resmsg[1]  = right Offset busy
this is the script that failed:
set_position
camera, EXP=30.0, FRAMES=1, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, ITER=1, IMAGE=STANDARD
pause
set_position
camera, EXP=30.0, FRAMES=1, COADDS=1, FILTER=4
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, ITER=1, IMAGE=STANDARD
  • preset again - change filter to J finish at elev 56.

12:05 preset at BS9135 (elev 35)
  • this script seems to be doing correctly the dither sequence
set_position
camera, EXP=40.0, FRAMES=1, COADDS=1, FILTER=4
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, ITER=1, IMAGE=STANDARD
pause
offset, COORD=detxy, X=-3.0, Y=3.0
camera, EXP=20.0, FRAMES=1, COADDS=1, FILTER=5
dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, ITER=1, IMAGE=STANDARD

12:30 AO people cannot do anything now, we try a 4th standard waiting for the sky flats
  • preset BS9126 (elev 78)

12:55 put dark filter and do some darks while we wait for flats
  • need to put /no_tcs so that it does not ask for the AO loop

13:14 observe, flatJ, /no_tcs * flats in H and in K BUT the counts are coming up too slow...

13:41 we cannot wait for high K flats because we need to hand over the telescope
  • tomorrow we begin early in the afternoon with flats

13:47 observe, darkCOR, /no_tcs
  • we do only 5frames of 600sec so that we can calculate corquad coefficients. During the first exposure time the lights in the dome were turn on. And Pisces is filling up. Possible the darks are not good

Quicklook

Test of the quicklook showing problems when multiple IDL license running.

comments on observe routines

We are building our experience using the observing routine "Observe". We have seeing a bit confuse the way to do a preacquire image and to move the star in the position that we want in the array. It will be interesting to have the commands to acquire images and to send offsets from the observe without using the iif_ commands. -The filter wheel is not in full operations and when we set a filter change inside a script, this doen't move the filter(something that we can see from the webcam) and the scripts goes forward. -Inside a dither pattern every few seconds during the exposure time (integrating an image)the Observe requests to stop or to keep going the script (small window pop up). -Setting the default dither pattern, eg: dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, InformationTechnologyER=1, IMAGE=STANDARD the dither finished in the last position of the dither and probably it will be better to finish in the zero position. -A dither script where the turn back the star to the initial position was:

set_position camera, EXP=10.0, FRAMES=1, COADDS=1, FILTER=4 dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, InformationTechnologyER=1, IMAGE=STANDARD pause offset, COORD=detxy, X=-3.0, Y=3.0 camera, EXP=20.0, FRAMES=1, COADDS=1, FILTER=5 dither, COORD=detxy, TYPE=box, SIZE=2, DIST=3.0, InformationTechnologyER=1, IMAGE=STANDARD

- It could be possible that the command set_position can be set as a single command, instead of request of iif_offset.

- the images obtained by observe are saved twice: once in the correct directory created by the *obs and once more in the main UTdirectory or in the scripts directory. There is also some confusion with the verbose mode... for example in yesterdays flat sequence in the IDL verbose we would see the first file twice and in the end it would scip the last image... in the directory though we found all images with the correct numbers...

-- KonstantinaBoutsia - 13 Oct 2011
Topic revision: r23 - 25 Jul 2014, DouglasFisher
This site is powered by FoswikiCopyright © by the contributing authors. All material on this collaboration platform is the property of the contributing authors.
Ideas, requests, problems regarding Foswiki? Send feedback